OpenStudy (anonymous):

find the coordinates of the focus of the parabola x^2=8(y-1)

OpenStudy (anonymous):

vertical parabola x^2=4py => 4p = 8 => p=2 the focus (0, 2)

OpenStudy (anonymous):

x^2=4py and u dont have to care about y-1

OpenStudy (anonymous):

hw come ?

OpenStudy (anonymous):

we have one more equation of vertical parabola: (x-h)^2 = 4p (y-k) with this u can see h=0 and k=1, so the center is ( 0,1). It doesnt affect to p: coordinate of focus

OpenStudy (anonymous):

the coordinates are wrong, the answers says (0,3)

OpenStudy (anonymous):

so heres the answer: (x-h)^2 = 4p (y-k) with this u can see h=0 and k=1, so the center is ( 0,1). so the focus would be (h, k+p) h=0 k+p = 1+2 = 3 my bad :)

OpenStudy (anonymous):

and the focus is on the axis of symmetry